### v3

@@ 4.1 Experimental Setting
  • MMLU: This is a collection of data examples that are in the Math category from the Massive Multitask Language Understanding (MMLU) benchmark (Hendrycks et al., 2020), focusing on numerical reasoning, symbolic manipulation, and procedural problem solving.
  • Metaphor (Tong et al., 2024): This dataset focuses on evaluating whether a highlighted word in context is used metaphorically in the context.
  • TruthfulQA (Lin et al., 2021): This dataset tests LLM trustworthy generation by posing questions with common misconceptions or false premises.
  • LogiQA (Liu et al., 2020): This dataset is designed for multi-step logical reasoning based on Chinese civil service exam questions.
Figure 3: The distribution of question length per dataset.
@@
  • CoT (Chain-of-Thought) (Wei et al., 2022): Prompts the model to think step-by-step for reasoning.
@@
  • ToT (Tree-of-Thought) (Yao et al., 2023): Structures reasoning path as a tree, exploring and selecting among multiple paths.
  • Best-of-N (Ji et al., 2023): Produces $N$ candidate chains, selects the best based on a predefined scoring metric.
  • Auto-CoT (Zhang et al., 2022): Auto-constructs reasoning chains by generated prompts, reducing manual work.
  • In-context CoT (ICL) (Brown et al., 2020): Leverages in-context CoT generation by presenting examples of few-shot train set directly within the prompt.
Evaluation and other details. To evaluate the alignment between LLM-generated responses and the ground truth, we adopt the “LLM-as-a-Judge" paradigm (Zheng et al., 2023), utilizing GPT-4o to assess both the semantic equivalence of answers and the quality of their explanations through dedicated judgment prompts, as illustrated in Figure 8. In each evaluation, the top_p parameter is set to 0.1 and the max_token parameter is set to 5,000, with no system prompt utilized. We random select 100 out of 1,000 samples as few-shot examples for AdaReasoner and ICL. ToT uses a beam width of 2 and a max length of 3. Baselines follow default settings with in-context examples from the same dataset and type. AdaReasoner uses a fixed learning rate of 0.01, BERT embeddings (768-d) for the input question, and a 3-layer MLP for each policy head.
@@ 4.2 Main Results
Figure 4 shows that when AdaReasoner is trained in few-shot scenarios, its performance exhibits marginal gains beyond 100 shots, universally for Qwen-2.5-72B, LLaMA-3.3-70B and GPT-4o. With 50–100 demonstrations suffice for the AdaReasoner to learn core reasoning patterns, validating the efficiency of the few-shot setting. This also match the theoretical result of fast convergence of the adapter supported in Appendix B.
@@ 4.3 Ablation Studies
To evaluate the effectiveness of Boltzmann exploration, we replace it by applying Thompson sampling (Russo et al., 2018) to all actions (w/ Bandit Adapter), which leads to a performance drop to 75.89%. To evaluate the effectiveness of the reward model, we added Gaussian noise ($\sigma=0.01$) to reward signal (w/ Perturbed Reward), and rescaled reward value from the interval [0 1] to the interval [-0.5 0.5] (w/ [-0.5 0.5] Reward). The results show that Adareasoner is robust to reward noise yet sophisticated in reward rescaling.
@@ 4.4 OOD Generalization of AdaReasoner
Table 3 shows if the AdaReasoner trained on the above-mentioned four datasets can be effectively applied on other out of domain (OOD) applications, such as multilingual emotion analysis BRIGHTER dataset (Muhammad et al., 2025), spatial planning in the StepGame dataset (Shi et al., 2022), and commonsense reasoning in the CRoW dataset (Ismayilzada et al., 2023). On the 150 QA pairs randomly sampled from each of these datasets that AdaReasoner has never encountered before, we can observe a stable superior performance of Adareasoner over other reasoning methods.
@@ 4.5 AdaReasoner on Knowledge Intensive Datasets
We next challenge our method on knowledge-intensive datasets, such as GPQA (Rein et al., 2024), MMLUChem (Hendrycks et al., 2020), and MedExQA (Kim et al., 2024), which require general domain knowledge or domain-specific knowledge in areas like chemistry, medicine. We randomly select 100 samples from each of these three datasets for training, and 500 samples for testing. As shown in Figure 5, AdaReasoner shows a modest yet consistent capacity to adjust to questions requiring intensive knowledge, outperforming conventional reasoning approaches such as CoT and ToT. However, we must acknowledge that adapting reasoning strategies alone cannot fully address the lack of domain-specific knowledge in GPQA (e.g., general facts, cultural references, history). A case-by-case analysis in Table 8 reveals that the adapter often selects self-audit, cross-reasoning, or creative prompt variants for such examples. Combining Table 8 with Table 7, the most frequently selected $a_{p}$ values—reflective self-questioning for logic-intensive tasks and creative assumption-challenging for Knowledge Intensive and Metaphor—suggest that cognitive configuration adaptation is a promising direction for further exploration, and this is just one of many intriguing patterns uncovered.
@@ Appendix B Theoretical Analysis of AdaReasoner
To support the empirical observations regarding AdaReasoner’s few-shot adaptation and robust performance across tasks, we present a theoretical analysis that characterizes its optimization bound and regret guarantees. We first analyze the error bound, and under the SGD condition, AdaReasoner achieves the $\frac{2\bigl{(}J(\Theta^{*})-J(\Theta_{0})\bigr{)}}{\eta\,K}\;+\;L\,\eta\,\sigma^{2}$ error bound. We then derive a regret bound for AdaReasoner’s softmax-based exploration policy using results from the non-stochastic multiarmed bandit theorem [Sutton and Barto, 2018]. This regret bound is provably sublinear, scaling as $O(\sqrt{K|\mathcal{A}|\log|\mathcal{A}|})$. Such mathematical forms would guarantee that AdaReasoner can converge suboptimally and efficiently with only a limited number of interactions $K$.
@@
To analyze the convergence of AdaReasoner in optimizing $\Theta$, we define the expected-reward objective as Equation 7:
@@
where $\beta=1/\tau$ is the inverse temperature of Boltzmann exploration  [Sutton and Barto, 2018].
@@
By viewing Softmax exploration as an instance of the exponential-weighting scheme, we can apply classical results from the non-stochastic multi-armed bandit problem, which yield the following bound for appropriately chosen $\beta$ [Auer et al., 2002]:
@@
Consequently, the per-step regret satisfies
@@
which vanishes rapidly as $K$ grows. In particular, once $K\gg|\mathcal{A}|\ln|\mathcal{A}|$, the average regret is negligible. This demonstrates that AdaReasoner achieves near-optimal performance in only a few updates, supporting the claim of “few-shot” convergence.
@@
Pareto frontier analysis based on regret
In Appendix B we showed that after $K$ training steps AdaReasoner’s cumulative regret satisfies
$R(K)\;=\;\sum_{k=1}^{K}\bigl{[}\mu_{k}(a^{*}_{k})-\mu_{k}(a_{k})\bigr{]}\;\leq\;O\!\bigl{(}\sqrt{K\,|\mathcal{A}|\,\ln|\mathcal{A}|}\bigr{)},$
so that
$\sum_{i=1}^{K}\bigl{[}f_{i}(\Theta^{*})-f_{i}(\Theta_{K})\bigr{]}\;=\;R(K),$
where $\Theta_{K}$ is the policy after $K$ steps and $\Theta^{*}$ the (hypothetical) joint-optimal policy. We now show this bound implies $\Theta_{K}$ lies on the Pareto frontier of the performance vector $\bigl{(}f_{1},\dots,f_{K}\bigr{)}$.
Theorem 2.
Let $\Theta_{K}$ be the policy learned after $K$ steps, and let
$f_{i}(\Theta)\;=\;\mathbb{E}_{a\sim\pi_{\Theta}(\cdot\mid q_{i})}\bigl{[}r(q_{i},a)\bigr{]}.$
Then for any other policy $\Theta^{\prime}$,
$\max_{i}\bigl{[}f_{i}(\Theta^{\prime})-f_{i}(\Theta_{K})\bigr{]}\;\leq\;R(K).$
In particular, if $K\gg|\mathcal{A}|\ln|\mathcal{A}|$ so that $R(K)$ is negligible, no $\Theta^{\prime}$ can strictly dominate $\Theta_{K}$, and hence $\Theta_{K}$ lies on the Pareto frontier.
Proof.
Suppose by contradiction that some $\Theta^{\prime}$ strictly dominates $\Theta_{K}$ by more than $R(K)$ on at least one dimension:
$\forall\,i:\;f_{i}(\Theta^{\prime})\geq f_{i}(\Theta_{K}),\quad\exists\,j:\;f_{j}(\Theta^{\prime})-f_{j}(\Theta_{K})\;>\;R(K).$
Summing over $i=1,\dots,K$ yields
$J(\Theta^{\prime})-J(\Theta_{K})=\sum_{i=1}^{K}\bigl{[}f_{i}(\Theta^{\prime})-f_{i}(\Theta_{K})\bigr{]}\;>\;R(K).$
On the other hand, from the regret bound we have
$J(\Theta^{*})-J(\Theta_{K})=\sum_{i=1}^{K}\bigl{[}f_{i}(\Theta^{*})-f_{i}(\Theta_{K})\bigr{]}=R(K),$
and since $J(\Theta^{\prime})\leq J(\Theta^{*})$ by definition of the optimal policy,
$J(\Theta^{\prime})-J(\Theta_{K})\;\leq\;R(K),$
a contradiction. Therefore no $\Theta^{\prime}$ can improve any $f_{i}$ by more than $R(K)$ without sacrificing another dimension, and $\Theta_{K}$ is on the Pareto frontier when $R(K)$ is negligible. ∎
@@ Appendix D Dataset Examples
MMLU (Math) If $193^{x-3}+10=$ 19, then $x=$ (A) 3 (B) 4 (C) 5 (D) 6 5
@@ Appendix F Prompt Templates
The prompt templates adopted in this study are provided in Figure 8, Figure 9, and Figure 10. Figure 8 depicts the prompt format designed for binary judgment-based evaluation of LLM simulations. Figure 9 shows the template applied by AdaReasoner for generating responses. Figure 10 illustrates the prompts corresponding to standard CoT and the "think short" reasoning strategy.
Prompt Template Assess with rigorous precision whether the provided reasoning process matches the ground truth answer. For a given option and response, you need to match the content of the option and response. You must not rely on the option index only, as in many cases, the index is actually incorrect.
Apply these criteria for judgment and carefully consider: Mandatory Evaluation Criteria 1. Content Equivalence: Accept only fully equivalent numerical representations (e.g., 0.5, 50%, 1/2) and variations in units or notation when they completely match the ground truth. 2. Logical Inference: Verify that at least one reasoning step directly and logically deduces the entire correct answer in a mathematically or logically sound manner. 3. Substantive Matching: For multiple-choice questions, assess the complete content of the answer (e.g., ensure "Option B" is fully equivalent to the correct answer, not just matching the label). 4. Semantic and Methodological Equivalence: Recognize alternative phrasing or solution methods only if a single step unambiguously converges on the complete correct answer. 5. Scientific and Technical Rigor: In technical contexts, differences in terminology, notation, or intermediate steps are acceptable only when they lead clearly and entirely to the correct conclusion. Using the criteria outlined above, determine whether any single rule is met–if so, the response is considered a match.
Question
@@
2. Question: {question}
3. Adaptive Reasoning Strategy
@@
You must follow no more than {optimal_steps} reasoning steps.
Requirements: 1. Provide one answer that completely satisfies the question’s requirements. 2. Ensure your reasoning strictly adheres to the specified steps and covers all necessary details. 3. Deliver a clear, precise, and accurate answer. 4. Avoid repetition or ambiguity; your response should be distinct and well-reasoned.
@@
Prompt Template Please think step by step to solve the question. / Please respond fastt and think quick when solving the question.
Question: {question}
Requirements: 1. Provide one answer that completely satisfies the question’s requirements. 2. Ensure your reasoning strictly adheres to the specified steps and covers all necessary details. 3. Deliver a clear, precise, and accurate answer. 4. Avoid repetition or ambiguity; your response should be distinct and well-reasoned.

### v4

@@ 4.1 Experimental Setting
  • MMLU: This is a collection of data examples that are in the Math category from the Massive Multitask Language Understanding (MMLU) benchmark [12], focusing on numerical reasoning, symbolic manipulation, and procedural problem solving.
  • Metaphor [49]: This dataset focuses on evaluating whether a highlighted word in context is used metaphorically in the context.
  • TruthfulQA [21]: This dataset tests LLM trustworthy generation by posing questions with common misconceptions or false premises.
  • LogiQA [23]: This dataset is designed for multi-step logical reasoning based on Chinese civil service exam questions.
@@
  • CoT (Chain-of-Thought) [54]: Prompts the model to think step-by-step for reasoning.
@@
  • ToT (Tree-of-Thought) [57]: Structures reasoning path as a tree, exploring and selecting among multiple paths.
  • Best-of-N [16]: Produces $N$ candidate chains, selects the best based on a predefined scoring metric.
  • Auto-CoT [58]: For each query, retrieve semantically nearest exemplars from a few-shot pool (via embedding clustering), generate CoT rationales, and concatenate the question–rationale–answer triplets as the in-context prompt; other settings follow the original.
  • In-context CoT (ICL) [5]: Leverages in-context CoT generation by presenting examples of few-shot train set directly within the prompt.
Evaluation and other details. To evaluate the alignment between LLM-generated responses and the ground truth, we adopt the “LLM-as-a-Judge” paradigm [60], utilizing GPT-4o to assess both the semantic equivalence of answers and the quality of their explanations through dedicated judgment prompts, as illustrated in Figure 8. In each evaluation, the top_p parameter is set to 0.1 and the max_token parameter is set to 5,000, with no system prompt utilized. We random select 100 out of 1,000 samples as few-shot examples for AdaReasoner and ICL. ToT uses a beam width of 2 and a max length of 3. Baselines follow default settings with in-context examples from the same dataset and type. AdaReasoner uses a fixed learning rate of 0.01, BERT embeddings (768-d) for the input question, and a 3-layer MLP for each policy head.
@@ 4.2 Main Results
In addition, to further demonstrate the reliability of LLM-as-Judge method used, we provide human annotated result in Appendix  H.
@@
Figure 3 shows that when AdaReasoner is trained in few-shot scenarios, its performance exhibits marginal gains beyond 100 shots, universally for Qwen-2.5-72B, LLaMA-3.3-70B and GPT-4o. With 50–100 demonstrations suffice for the AdaReasoner to learn core reasoning patterns, validating the efficiency of the few-shot setting. Theoretical worst case regret convergence as in Appendix B is $\mathcal{O}\sqrt{\mathcal{|A|}\text{ln}\mathcal{|A|}}$, AdaReasoner converges far faster in practice. This might be due to a shared encoder enforcing $\beta$-smoothness and near-convexity [8], and a pretrained reward model providing high-fidelity, high-SNR feedback (“warm start”) that accelerates few-shot policy updates.
@@ 4.3 Ablation Studies
To evaluate the effectiveness of Boltzmann exploration, we replace it by applying Thompson sampling [39] to all actions (w/ Bandit Adapter), which leads to a performance drop to 75.89%. To evaluate the effectiveness of the reward model, we added Gaussian noise ($\sigma=0.01$) to reward signal (w/ Perturbed Reward), and rescaled reward value from the interval [0 1] to the interval [-0.5 0.5] (w/ [-0.5 0.5] Reward). The results show that Adareasoner is robust to reward noise yet sophisticated in reward rescaling.
Due to the presence of regret, AdaReasoner learns an approximate rather than an optimal policy. To assess this, we analyze perturbed variants (Close and Distant Perturb) by selecting neighboring actions in embedding space by similarity. We also evaluate an Ensemble setting that aggregates independently trained policy heads without shared layers, further validating AdaReasoner’s design.
@@
Adareasoner (Close-perturb) 66.05 79.39 85.18 80.03 77.66
Adareasoner (Distant-perturb) 57.69 71.77 81.42 74.96 71.46
Adareasoner (Emsemble) 65.73 79.54 84.71 80.04 77.50
@@ 4.4 OOD Generalization of AdaReasoner
Table 3 shows if the AdaReasoner trained on the above-mentioned four datasets can be effectively applied on other out of domain (OOD) applications, such as multilingual emotion analysis BRIGHTER dataset [29], spatial planning in the StepGame dataset [40], and commonsense reasoning in the CRoW dataset [14]. On the 150 QA pairs randomly sampled from each of these datasets that AdaReasoner has never encountered before, we can observe a stable superior performance of Adareasoner over other reasoning methods.
@@ 4.5 AdaReasoner on Knowledge Intensive Datasets
We next challenge our method on knowledge-intensive datasets, such as GPQA [34], MMLUChem [12], and MedExQA [18], which require general domain knowledge or domain-specific knowledge in areas like chemistry, medicine. We randomly select 100 samples from each of these three datasets for training, and 500 samples for testing. As shown in Figure 4, AdaReasoner shows a modest yet consistent capacity to adjust to questions requiring intensive knowledge, outperforming conventional reasoning approaches such as CoT and ToT. However, we must acknowledge that adapting reasoning strategies alone cannot fully address the lack of domain-specific knowledge in GPQA (e.g., general facts, cultural references, history). A case-by-case analysis in Table 8 reveals that the adapter often selects self-audit, cross-reasoning, or creative prompt variants for such examples. Combining Table 8 with Table 7, the most frequently selected $a_{p}$ values—reflective self-questioning for logic-intensive tasks and creative assumption-challenging for Knowledge Intensive and Metaphor—suggest that cognitive configuration adaptation is a promising direction for further exploration, and this is just one of many intriguing patterns uncovered.
@@ Appendix B Theoretical Analysis of AdaReasoner
To support the empirical observations regarding AdaReasoner’s few-shot adaptation and robust performance across tasks, we present a theoretical analysis that characterizes its optimization bound and regret guarantees. We first analyze the error bound, and under the SGD condition, AdaReasoner achieves the $\frac{2\bigl(J(\Theta^{*})-J(\Theta_{0})\bigr)}{\eta\,K}\;+\;L\,\eta\,\sigma^{2}$ error bound. We then derive a regret bound for AdaReasoner’s softmax-based exploration policy using results from the non-stochastic multiarmed bandit theorem [47]. This regret bound is provably sublinear, scaling as $O(\sqrt{K|\mathcal{A}|\log|\mathcal{A}|})$. Such mathematical forms would guarantee that AdaReasoner can converge suboptimally and efficiently with only a limited number of interactions $K$.
@@
To analyze the convergence of AdaReasoner in optimizing $\Theta$, we define the expected‐reward objective as Equation 7:
@@
where $\beta=1/\tau$ is the inverse temperature of Boltzmann exploration  [47].
@@
By viewing Softmax exploration as an instance of the exponential‐weighting scheme, we can apply classical results from the non-stochastic multi-armed bandit problem, which yield the following bound for appropriately chosen $\beta$ [3]:
@@
Consequently, the per‐step regret satisfies
@@
which vanishes rapidly as $K$ grows. In particular, once $K\gg|\mathcal{A}|\ln|\mathcal{A}|$, the average regret is negligible. This demonstrates that AdaReasoner achieves near‐optimal performance in only a few updates, supporting the claim of “few-shot” convergence.
@@ Appendix D Dataset Examples
MMLU (Math) If $3^{x-3}+10=19$, then $x=$ (A) 3 (B) 4 (C) 5 (D) 6 5
@@ Appendix F Prompt Templates
The prompt templates adopted in this study are provided in Figure 8, Figure 9, and Figure 10. Figure 8 depicts the prompt format designed for binary judgment-based evaluation of LLM simulations. Figure 9 shows the template applied by AdaReasoner for generating responses. Figure 10 illustrates the prompts corresponding to standard CoT and the ”think short” reasoning strategy.
Prompt Template Assess with rigorous precision whether the provided reasoning process matches the ground truth answer. For a given option and response, you need to match the content of the option and response. You must not rely on the option index only, as in many cases, the index is actually incorrect. Apply these criteria for judgment and carefully consider: Mandatory Evaluation Criteria 1. Content Equivalence: Accept only fully equivalent numerical representations (e.g., 0.5, 50%, 1/2) and variations in units or notation when they completely match the ground truth. 2. Logical Inference: Verify that at least one reasoning step directly and logically deduces the entire correct answer in a mathematically or logically sound manner. 3. Substantive Matching: For multiple-choice questions, assess the complete content of the answer (e.g., ensure ”Option B” is fully equivalent to the correct answer, not just matching the label). 4. Semantic and Methodological Equivalence: Recognize alternative phrasing or solution methods only if a single step unambiguously converges on the complete correct answer. 5. Scientific and Technical Rigor: In technical contexts, differences in terminology, notation, or intermediate steps are acceptable only when they lead clearly and entirely to the correct conclusion. Using the criteria outlined above, determine whether any single rule is met–if so, the response is considered a match. Question
@@
2. Question: {question} 3. Adaptive Reasoning Strategy
@@
You must follow no more than {optimal_steps} reasoning steps. Requirements: 1. Provide one answer that completely satisfies the question’s requirements. 2. Ensure your reasoning strictly adheres to the specified steps and covers all necessary details. 3. Deliver a clear, precise, and accurate answer. 4. Avoid repetition or ambiguity; your response should be distinct and well-reasoned.
@@
Prompt Template Please think step by step to solve the question. / Please respond fastt and think quick when solving the question. Question: {question} Requirements: 1. Provide one answer that completely satisfies the question’s requirements. 2. Ensure your reasoning strictly adheres to the specified steps and covers all necessary details. 3. Deliver a clear, precise, and accurate answer. 4. Avoid repetition or ambiguity; your response should be distinct and well-reasoned.